You borrow $30,000 at an annual rate of 7% and must repay it in 7 equal installments at the end of each of the next 7 years. How much will you still owe at the end of the second year, after you make the second payment?
Answer #1

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

30,000=Annuity[1-(1.07)^-7]/0.07

30,000=Annuity*5.389289402

Annuity=30,000/5.389289402

=$5566.60(Approx)

Interest payment for 1st year=$30,000*7%=$2100

Hence principal payment for 1st year=(5566.6-2100)=$3466.6

Balance due after 1st year=(30,000-3466.6)=$26533.4

Interest payment for second year=$26533.4*7%=$1857.338

Hence principal payment for 2nd year =$5566.60-1857.338=$3709.262

Hence owed balance after 2nd year=26533.4-3709.262

=$22824.14(Approx).